I have a quick question for my Organic Chemistry class about the topic of alkene addition reactions. Its specifically about oxidative cleavage of alkenes. Can there only be a cleavage in a double bond only if the reagents of the reaction contains O3 (ozone)?
Answer #1

Ozonolysis is an organic reaction where the double bonds of alkene are cleaved with ozone. Alkenes can be oxidised with O3 to form alcohols , ketones or aldehydes or carboxylic acid .

No , there not only cleavage in a double bond only , but C-H bond oxidises to C-OH .

example - CH3-C(CH3)=CH-CH3(O3/ H2O2) CH3-C=O-CH3 CH3-COOH  

ketone carboxylic acid